OpenStudy (anonymous):

CALCULUS: how to do this without L'Hospital's rule lim x->2 (sqrt(6-x)-2)/(sqrt(3-x)-1)

OpenStudy (anonymous):

multiply by the conjugate. maybe twice

OpenStudy (anonymous):

@satellite73 i did but it gave me this and im stuck this seems too complicated to resolve.. http://puu.sh/14T7n

OpenStudy (anonymous):

plz help anyone? :S

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

anyone still on this question? o_O

OpenStudy (zarkon):

\[\lim_{x\to 2}\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1}\] \[=\lim_{x\to 2}\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1}\frac{\sqrt{6-x}+2}{\sqrt{6-x}+2}\frac{\sqrt{3-x}+1}{\sqrt{3-x}+1}\]

OpenStudy (zarkon):

simplify...take limit

OpenStudy (anonymous):

ok i just did, and it returned to its original form...

OpenStudy (zarkon):

\[=\lim_{x\to 2}\frac{6-x-4}{3-x-1}\frac{1}{\sqrt{6-x}+2}\frac{\sqrt{3-x}+1}{1}\]

OpenStudy (anonymous):

ya you see, it returns to the original form..

OpenStudy (anonymous):

http://puu.sh/14TOp

OpenStudy (zarkon):

\[=\lim_{x\to 2}\frac{2-x}{2-x}\frac{1}{\sqrt{6-x}+2}\frac{\sqrt{3-x}+1}{1}\] \[=\lim_{x\to 2}\frac{\cancel{2-x}}{\cancel{2-x}}\frac{1}{\sqrt{6-x}+2}\frac{\sqrt{3-x}+1}{1}\] \[=\lim_{x\to 2}\frac{\sqrt{3-x}+1}{\sqrt{6-x}+2}\]

OpenStudy (anonymous):

OMFG MY BAD

OpenStudy (anonymous):

IT CAN ACTUALLY BE SOLVED NOW!! THANKS!!!!

OpenStudy (zarkon):

please don't write 'OMFG'